# Unifying Feature-Based Explanations with Functional ANOVA and Cooperative Game Theory

### v1

@@ 3.1 Influence of Feature Distributions
We now illustrate differences between the fANOVA decompositions based on features from a two-dimensional multivariate normal distribution $X\sim\mathcal{N}(\mu,\Sigma)$ with mean $\mu\in\mathbb{R}^{2}$ and covariance $\Sigma\in\mathbb{R}^{2\times 2}$. We consider a linear function $F_{\text{lin}}(x)=\beta_{1}x_{1}+\beta_{2}x_{2}$, and one including a two-way interaction $F_{\text{2int}}(x)=F_{\text{lin}}(x)+\beta_{12}x_{1}x_{2}$. The main and interaction effects of the fANOVA decompositions are summarized in Table 1. For b- and m-fANOVA of $F_{\text{lin}}$ the main effect is intuitively the deviation of $x_{i}$ from $b_{i}$ and the expectation of $\mathbb{E}[X_{i}]=\mu_{i}$, respectively, which is scaled by the linear coefficient $\beta_{i}$ (first row). For c-fANOVA, this effect is distorted by the cross-correlation with the second feature. As stated in Theorem 2, it is clear that c-fANOVA reduces to m-fANOVA for independent features and that m-fANOVA equals b-fANOVA for $b_{i}:=\mathbbm{E}[X_{i}]$.
For $F_{2\text{int}}$, the first order effects of b- and m-fANOVA again only depend on feature $i$, but also include the interaction coefficient $\beta_{ij}$ (second row). In contrast, c-fANOVA additionally includes cross-correlation effects based on effects of features correlated with $x_{i}$. The two-way interaction of b- and m-fANOVA is solely defined by the centered interaction effect between the two features. Yet, c-fANOVA accounts again for cross-correlation effects. For a more detailed discussion, we refer to Section D.1.
@@ 3.2 Explanation Games via fANOVA
Corollary 1 follows directly from the definitions of the MT and the fANOVA components. In fact, both are special cases of the Möbius inversion theorem (Rota,, 1964, Proposition 2) with the inclusion ordering $\subseteq$” . While the MT is based on $F_{S}(x_{0})$, the fANOVA effect is based on the functions $F_{S}$.
@@
For independent features, the MT of the sensitivity game is the variance of the fANOVA effect $f_{S}$. For dependent features, the covariances are exclusively distributed among the MT. Lastly, for the risk game the MT $m^{(\text{risk})}(S)$ is the additive improvement in risk obtained by $S$. Given an explanation game, we now specify different types of feature influence.
@@ 3.4 Quantifying Feature Influence
We now illustrate the influence of main and higher-order effects for different feature influence methods. We consider three independent features $X_{i}\sim\mathcal{N}(0,1)$ and a functional relationship of $F_{3\text{int}}(x)=x_{1}+x_{2}+x_{3}+x_{1}x_{2}+x_{1}x_{2}x_{3}$. In this case, c-fANOVA is equivalent to m-fANOVA and b-fANOVA with $b=\mathbbm{E}[X]$, and each term in $F$ corresponds to a fANOVA effect. The weights of fANOVA effects for different feature influences of the local explanation are summarized in Table 3. For individual contributions (rows 1-3), the pure effect captures the main effect $x_{1}$, whereas the partial effect includes the equally distributed interactions $x_{1}x_{2}$ and $x_{1}x_{2}x_{3}$. Moreover, the full effect $\phi^{+}$ accounts for both interactions with weight $1$. For joint influence (rows 4-6), the pure effect captures the interaction $x_{1}x_{2}$, as well as the individual effects $x_{1},x_{2}$. The partial and full effects additionally capture the higher-order interaction $x_{1}x_{2}x_{3}$. Lastly, for interactions (rows 7-9), the pure effect captures the interaction $x_{1}x_{2}$, whereas the partial and full effect also account for the higher-order interaction $x_{1}x_{2}x_{3}$.
@@ 4.3 A Guide for Practitioners
Computational complexity is typically assessed by the number of model calls to $F$. In the local explanation game the primary factor is the value function $F_{S}$, for which a single model call suffices for b-fANOVA, whereas m- and c-fANOVA rely on approximations of expectations, requiring varying numbers of evaluations depending on the desired estimation quality (Baniecki et al.,, 2024). For global explanations, local explanation game values must be computed across the feature distribution or the labeled data. Once game values are available, pure and full individual and joint effects each require two game evaluations, while pure and full interactions for a set of size $s$ demand $2^{s}$ evaluations. Partial effects are the most computationally expensive, necessitating $2^{d}$ game evaluations or their approximations. Importantly, game evaluations can be reused across different explanation games, types, and effects.
@@ 5 EXPERIMENTS
We now showcase our framework on synthetic and four real-world datasets. For m-fANOVA, we use a background dataset of $512$ samples, and for c-fANOVA the synthetic ground-truth. Further details and experiments are described in Section F.3 and Appendix G. Our code is available at https://github.com/FFmgll/unifying_feature_based_explanations.
@@ A.2 Proof of Theorem 2
Above result for multilinear functions can be extended to linear combinations of multilinear functions, since expectations are linear. In summary, every model $F$ that can be described as the sum of several weighted multilinear functions will have similar b- and m-fANOVA effects for independent features, if $b:=\mathbb{E}[X]$.
@@ B.2 From Decompositions to Influence Measures and Back
Using a suitable type of imputation method, and thereby definition of $F_{S}$, we obtain a fANOVA decomposition. This ensured by the recursive definition of fANOVA effects from Eq. (1). In fact, for the last effect, we have
@@
which yields a decomposition, as described for instance by Apley and Zhu, (2020); Iwasawa and Matsumori, (2024). Moreover, this recursive definition related to the MT also ensures the orthogonality-like property (Apley and Zhu,, 2020; Iwasawa and Matsumori,, 2024), since the MT of a game with a single MT is zero, unless these are the same. Notably, partial individual, joint and interaction effects summarize this fANOVA again into a functional decomposition (Bordt and von Luxburg,, 2023) with components up to order $k$, i.e.
@@ C.2 Background on Cooperative Game Theory
In this section, we briefly cover the definitions of GVs (Marichal et al.,, 2007) and SIs Interaction indices (II) (Fujimoto et al.,, 2006). Both were introduced based on axioms as a weighted sum of joint marginal contributions and discrete derivatives as
@@ D.1.1 Baseline fANOVA (b-fANOVA)
$=\beta_{1}x_{1}+\beta_{2}x_{2}-\beta_{1}(x_{1}-b_{1})-\beta_{2}(x_{2}-b_{2})-(\beta_{1}b_{1}+\beta_{2}b_{2}))$
@@
It follows that $F_{\text{lin}}(x)=f^{(b)}_{\emptyset}+f^{(b)}_{1}(x)+f^{(b)}_{2}(x)$, hence, the function can be decomposed into a constant plus the first-order centered (main) effects of the two features. The centering is based on the baseline value $b$ for b-fANOVA.
@@ D.1.2 Marginal fANOVA (m-fANOVA)
It follows that $F_{\text{lin}}(x)=f^{(m)}_{\emptyset}+f^{(m)}_{1}(x)+f^{(m)}_{2}(x)$, hence, the function can be decomposed into a constant plus the first-order mean-centered (main) effects of the two features. As shown in Theorem 2, b-fANOVA corresponds to m-fANOVA for this linear functional relationship if $b:=\mu$.
@@
$=\beta_{1}x_{1}+\beta_{2}x_{2}+\beta_{12}x_{1}x_{2}-(\beta_{1}(x_{1}-\mu_{1})+\beta_{12}\mu_{2}(x_{1}-\mu_{1})-\beta_{12}\sigma_{12}))$
@@ D.1.3 Conditional fANOVA (c-fANOVA)
$-\beta_{2}(x_{2}-\mu_{2})-(x_{2}-\mu_{2})\frac{\beta_{1}\sigma_{12}}{\mathbb{V}[X_{2}]}-(\beta_{1}\mu_{1}+\beta_{2}\mu_{2}))$
$=-\sum_{\ell\in 12}(x_{\ell}-\mu_{\ell})\frac{\beta_{-\ell}\sigma_{-\ell,\ell}}{\mathbb{V}[X_{\ell}]}$
@@
$-(\beta_{i}\mu_{i}+\beta_{j}\mu_{j}+\beta_{ij}(\sigma_{ij}+\mu_{i}\mu_{j}))$
@@
$-\sum_{\ell\in 12}(x_{\ell}-\mu_{\ell})\frac{\beta_{-\ell}\sigma_{-\ell,\ell}}{\mathbb{V}[X_{\ell}]}-\beta_{12}\sigma_{12}\left(\frac{(x_{1}-\mu_{1})x_{1}}{\mathbb{V}[X_{1}]}+\frac{(x_{2}-\mu_{2})x_{2}}{\mathbb{V}[X_{2}]}\right)$
@@ D.2 Details on Illustration of Feature Influence Methods
In this section, we provide more detailed definitions of the feature influence methods of the illustrative example of Section 3.4. As defined in Section 3.4, we assume that we are given three independent and standard normally distributed features $X_{1},X_{2},X_{3}\sim\mathcal{N}(0,1)$ and the following functional relationship: $F(x)=x_{1}+x_{2}+x_{3}+x_{1}x_{2}+x_{1}x_{2}x_{3}$. In this case the m-fANOVA is equivalent to the c-fANOVA and the b-fANOVA with $b=\mathbbm{E}[X]$. We now define the feature influence measures based on the fANOVA decomposition. The respective weights of each effect are summarized in Table 3 in the main paper.
@@
Note that for a linear functional relationship without interactions and uncorrelated features, the individual influence methods result in the same contribution defined by the main effect: $\phi^{\emptyset}(i)=\phi^{SV}(i)=\phi^{+}(i)=\beta_{i}\cdot(x_{i}-\mu_{i})$, while the interaction influence methods become zero: $\phi^{I\emptyset}(S)=\phi^{SI}(S)=\phi^{I+}(S)=0$, and the joint influence methods are defined by the weighted sum of main effects $w_{1}(x_{S}-\mu_{S})\beta_{S}$, where $w_{1}=1$ for $\phi^{\emptyset}(S)$ and $\phi^{+}(S)$.
@@ E.1.1 Perturbation-Based Explanations
PredDiff-1 and PredDiff-Patch.
PredDiff (Zintgraf et al.,, 2017) measures similarly to Occlusion the change in prediction when masking features. However, instead of using a baseline, PredDiff uses the conditional distribution, i.e., for a patch $S\subseteq D$
@@
Hence, PredDiff-Patch measures the full joint effect, and PredDiff-1 the full individual effect of c-fANOVA.
@@ E.1.2 Gradient-Based Attribution Methods
DeepSHAP (Lundberg and Lee,, 2017) computes attributions layer-wise, where the SV is used. It was shown by Deng et al., (2024), that for a each layer, the layer-wise attribution is given by
$\phi(i)=\sum_{\kappa\in\Omega_{i}}\phi(\kappa)+\sum_{i\in S:s>1}\frac{1}{s}\sum_{\kappa\in\Omega_{S}}I(\kappa)=\psi(i)+\sum_{S\supset i}\frac{1}{s}J(S)=f_{i}^{(b)}+\sum_{S\supset i}\frac{1}{s}f^{(b)}_{S},$
@@
i.e., applying IG twice for each feature. By taking the gradient, all effects vanish that do not involve feature $x_{i}$, as seen in Eq. (5). In detail, for IG, we have
@@ F.1 Synthetic Examples
To model the data generating processes $F_{\text{lin}}$, $F_{\text{int}}$, and $F_{\text{add}}$, we randomly draw 10,000 samples and fit correctly specified linear models with sklearn (Pedregosa et al.,, 2011), ensuring correctly learned effects for each functional form. Input features are transformed to match the underlying functional relationships, with the relevant main effects and interaction terms selected accordingly. This method guarantees that the models retain linearity in the parameter space, consistent with the specified functional relationships. As a result, we obtain an almost perfect fit across all three synthetic datasets, achieving $R^{2}$ values greater than 0.99 on a separately sampled test dataset of size 10,000 following the same distribution as defined above.
@@ F.2 Models and Datasets for Real-World Examples
  • Random Forest: A random forest model with 50 trees was trained on the preprocessed data, achieving an $R^{2}=0.9471\pm 0.0024$ and a MSE =1 863±103absentplus-or-minus1863103=1\,863$\pm$ 103= 1 863 ± 103.
  • Neural Network: A fully connected neural network with four hidden layers (sizes 100, 50, 10, 1) was trained on the preprocessed data, where the categorical features were treated as numeric and standardized. The regression target variable was additionally log-normalized. This model achieved an $R^{2}=0.9436\pm 0.0016$ and an MSE of $0.3299\pm 0.0055$.
@@
The Titanic dataset is used for a binary classification task, predicting passenger survival based on features such as age, gender, class of travel, and fare. We pre-process the data by applying an OrdinalEncoder to the categorical features.
@@
The sentiment analysis language data is based on the IMDB dataset (Maas et al.,, 2011; Lhoest et al.,, 2021) where example sentences to the model are self-created. The model is a pre-trained and fine-tuned version of DistilBERT (Sanh et al.,, 2019) on IMDB as provided with the transformers (Wolf et al.,, 2020) framework. The model predicts the sentiment of movie reviews with a sentiment score $\left[-1,1\right]$ where values close to $-1$ denote a negative predicted sentiment and values close to $1$ correspond to a positive predicted sentiment. Sentiment scores close to 00 are considered neutral. In our experiments, we manually remove punctuation and special symbols. In case of b-fANOVA, we set out-of-coalition tokens to the MASK token of the transformer model. Notably, the selected DistilBERT variant assigns sentences provided only MASK tokens moderately high scores of around $0.50$ (i.e. on the positive side). We do not evaluate and report performance scores of the pre-trained model on a separate test set.

### v2

@@ 3.1 Influence of Feature Distributions
We now illustrate differences between the fANOVA decompositions based on features from a two-dimensional multivariate normal distribution $X\sim\mathcal{N}(\mu,\Sigma)$ with mean $\mu\in\mathbb{R}^{2}$ and covariance $\Sigma\in\mathbb{R}^{2\times 2}$. We consider a linear function $F_{\text{lin}}(x)=\beta_{1}x_{1}+\beta_{2}x_{2}$, and one including a two-way interaction $F_{\text{2int}}(x)=F_{\text{lin}}(x)+\beta_{12}x_{1}x_{2}$. The main and interaction effects of the fANOVA decompositions are summarized in Table 1. For b- and m-fANOVA of $F_{\text{lin}}$ the main effect is intuitively the deviation of $x_{i}$ from $b_{i}$ and the expectation of $\mathbb{E}[X_{i}]=\mu_{i}$, respectively, which is scaled by the linear coefficient $\beta_{i}$ (first row). For c-fANOVA, this effect is distorted by the cross-correlation with the second feature. As stated in Theorem 2, it is clear that c-fANOVA reduces to m-fANOVA for independent features and that m-fANOVA equals b-fANOVA for $b_{i}:=\mathbb{E}[X_{i}]$. For $F_{2\text{int}}$, the first order effects of b- and m-fANOVA again only depend on feature $i$, but also include the interaction coefficient $\beta_{ij}$ (second row). In contrast, c-fANOVA additionally includes cross-correlation effects based on effects of features correlated with $x_{i}$. The two-way interaction of b- and m-fANOVA is solely defined by the centered interaction effect between the two features and thus results in zero values for $F_{lin}$. Yet, c-fANOVA accounts again for cross-correlation effects. For a more detailed discussion, we refer to Section D.1.
@@ 3.2 Explanation Games via fANOVA
Corollary 1 follows directly from the definitions of the MT and the fANOVA components. In fact, both are special cases of the Möbius inversion theorem (Rota,, 1964, Proposition 2) with the inclusion ordering. While the MT $m^{\text{(loc)}}_{x_{0}}$ is based on $F_{S}(x_{0})$, the fANOVA effect is based on the functions $F_{S}$.
@@
For independent features, the MT of the sensitivity game is the variance of the fANOVA effect $f_{S}$. For dependent features, the covariances are exclusively distributed among the MT. Lastly, for the risk game, the MT $m^{(\text{risk})}(S)$ is the additive improvement in risk obtained by $S$. Given an explanation game, we now specify different types of feature influence.
@@ 3.4 Quantifying Feature Influence
We now illustrate the influence of main and higher-order effects for different feature influence methods. We consider three independent features $X_{i}\sim\mathcal{N}(0,1)$ and a functional relationship of $F_{3\text{int}}(x)=x_{1}+x_{2}+x_{3}+x_{1}x_{2}+x_{1}x_{2}x_{3}$. In this case, c-fANOVA is equivalent to m-fANOVA and b-fANOVA with $b=\mathbb{E}[X]$, and each term in $F$ corresponds to a fANOVA effect. The weights of fANOVA effects for different feature influences of the local explanation are summarized in Table 3. For individual contributions (rows 1-3), the pure effect captures the main effect $x_{1}$, whereas the partial effect includes the equally distributed interactions $x_{1}x_{2}$ and $x_{1}x_{2}x_{3}$. Moreover, the full effect $\phi^{+}$ accounts for both interactions with weight $1$. For joint influence (rows 4-6), the pure effect captures the interaction $x_{1}x_{2}$, as well as the individual effects $x_{1},x_{2}$. The partial and full effects additionally capture the higher-order interaction $x_{1}x_{2}x_{3}$. Lastly, for interactions (rows 7-9), the pure effect captures the interaction $x_{1}x_{2}$, whereas the partial and full effect also account for the higher-order interaction $x_{1}x_{2}x_{3}$.
@@ 4.3 A Guide for Practitioners
Computational complexity is typically assessed by the number of model calls to $F$. In the local explanation game the primary factor is the value function $F_{S}$, for which a single model call suffices for b-fANOVA, whereas m- and c-fANOVA rely on approximations of expectations, requiring varying numbers of evaluations depending on the desired estimation quality (Baniecki et al.,, 2025). For global explanations, local explanation game values must be computed across the feature distribution or the labeled data. Once game values are available, pure and full individual and joint effects each require two game evaluations, while pure and full interactions for a set of size $s$ demand $2^{s}$ evaluations. Partial effects are the most computationally expensive, necessitating $2^{d}$ game evaluations or their approximations. Importantly, game evaluations can be reused across different explanation games, types, and effects.
@@ 5 EXPERIMENTS
We now showcase our framework on synthetic and real-world datasets. For m-fANOVA, we use a background dataset of $512$ samples, and for c-fANOVA the synthetic ground-truth. Further details and experiments are described in Section F.3 and Appendix G. Our code is available at https://github.com/FFmgll/unifying_feature_based_explanations.
@@ A.2 Proof of Theorem 2
Results derived from Theorem 2 for multilinear functions can be extended to linear combinations of multilinear functions, since expectations are linear. In summary, every model $F$ that can be described as the sum of several weighted multilinear functions will have similar b- and m-fANOVA effects for independent features, if $b:=\mathbb{E}[X]$.
@@ B.2 From Decompositions to Influence Measures and Back
Using a suitable type of imputation method, and thereby definition of $F_{S}$, we obtain a fANOVA decomposition. This is ensured by the recursive definition of fANOVA effects from Eq. (1). In fact, for the last effect, we have
@@
which yields a decomposition, as described for instance by Apley and Zhu, (2020); Iwasawa and Matsumori, (2024). Moreover, this recursive definition related to the MT also ensures the orthogonality-like property (Apley and Zhu,, 2020; Iwasawa and Matsumori,, 2024). Notably, partial individual, joint and interaction effects summarize this fANOVA again into a functional decomposition (Bordt and von Luxburg,, 2023) with components up to order $k$, i.e.
@@ C.2 Background on Cooperative Game Theory
In this section, we briefly cover the definitions of GVs (Marichal et al.,, 2007) and Interaction Indices (II) (Fujimoto et al.,, 2006). Both were introduced based on axioms as a weighted sum of joint marginal contributions and discrete derivatives as
@@ D.1.1 Baseline fANOVA (b-fANOVA)
$=\beta_{1}x_{1}+\beta_{2}x_{2}-\beta_{1}(x_{1}-b_{1})-\beta_{2}(x_{2}-b_{2})-(\beta_{1}b_{1}+\beta_{2}b_{2})$
@@
It follows that $F_{\text{lin}}(x)=f^{(b)}_{\emptyset}+f^{(b)}_{1}(x)+f^{(b)}_{2}(x)$, hence, the function can be decomposed into a constant plus the centered first-order (main) effects of the two features. The centering is based on the baseline value $b$ for b-fANOVA.
@@ D.1.2 Marginal fANOVA (m-fANOVA)
It follows that $F_{\text{lin}}(x)=f^{(m)}_{\emptyset}+f^{(m)}_{1}(x)+f^{(m)}_{2}(x)$, hence, the function can be decomposed into a constant plus the mean-centered first-order (main) effects of the two features. As shown in Theorem 2, b-fANOVA corresponds to m-fANOVA for this linear functional relationship if $b:=\mu$.
@@
$=\beta_{1}x_{1}+\beta_{2}x_{2}+\beta_{12}x_{1}x_{2}-(\beta_{1}(x_{1}-\mu_{1})+\beta_{12}\mu_{2}(x_{1}-\mu_{1})-\beta_{12}\sigma_{12})$
@@ D.1.3 Conditional fANOVA (c-fANOVA)
$-\beta_{2}(x_{2}-\mu_{2})-(x_{2}-\mu_{2})\frac{\beta_{1}\sigma_{12}}{\mathbb{V}[X_{2}]}-(\beta_{1}\mu_{1}+\beta_{2}\mu_{2})$
=$=-\sum_{\ell\in\{1,2\}}(x_{\ell}-\mu_{\ell})\frac{\beta_{-\ell}\sigma_{-\ell,\ell}}{\mathbb{V}[X_{\ell}]}$
@@
$-(\beta_{i}\mu_{i}+\beta_{j}\mu_{j}+\beta_{ij}(\sigma_{ij}+\mu_{i}\mu_{j})$
@@
$-\sum_{\ell\in\{1,2\}}(x_{\ell}-\mu_{\ell})\frac{\beta_{-\ell}\sigma_{-\ell,\ell}}{\mathbb{V}[X_{\ell}]}-\beta_{12}\sigma_{12}\left(\frac{(x_{1}-\mu_{1})x_{1}}{\mathbb{V}[X_{1}]}+\frac{(x_{2}-\mu_{2})x_{2}}{\mathbb{V}[X_{2}]}\right)$
@@ D.2 Details on Illustration of Feature Influence Methods
In this section, we provide more detailed definitions of the feature influence methods of the illustrative example of Section 3.4. As defined in Section 3.4, we assume that we are given three independent and standard normally distributed features $X_{1},X_{2},X_{3}\sim\mathcal{N}(0,1)$ and the following functional relationship: $F(x)=x_{1}+x_{2}+x_{3}+x_{1}x_{2}+x_{1}x_{2}x_{3}$. In this case the m-fANOVA is equivalent to the c-fANOVA and the b-fANOVA with $b=\mathbb{E}[X]$. We now define the feature influence measures based on the fANOVA decomposition. The respective weights of each effect are summarized in Table 3 in the main paper.
@@
Note that for a linear functional relationship without interactions and uncorrelated features, the individual influence methods result in the same contribution defined by the main effect: $\phi^{\emptyset}(i)=\phi^{SV}(i)=\phi^{+}(i)=\beta_{i}\cdot(x_{i}-\mu_{i})$, while the interaction influence methods become zero: $\phi^{I\emptyset}(S)=\phi^{SI}(S)=\phi^{I+}(S)=0$.
Similarly to individual influence methods, the pure, partial, and full effects result in the same contribution defined by the sum of the main effects: $\phi^{\emptyset}(S)=\phi^{GV}(S)=\phi^{+}(S)=\sum_{i\in S}\phi^{\emptyset}(i)=\sum_{i\in S}\beta_{i}\cdot(x_{i}-\mu_{i})$.
@@ E.1.1 Perturbation-Based Explanations
PredDiff-1, PredDiff-Patch and PredDiff-Interaction
PredDiff (Robnik-Sikonja and Kononenko,, 2008) is a popular method for image classification tasks (Zintgraf et al.,, 2017). It measures similarly to Occlusion the change in prediction when masking features. However, instead of using a baseline, PredDiff uses the conditional distribution, i.e., for a patch $S\subseteq D$
@@
Hence, PredDiff-Patch measures the full joint effect, and PredDiff-1 the full individual effect of c-fANOVA. PredDiff-Interaction (Blücher et al.,, 2022) measures the interaction by extracting individual effects, thereby measuring full interaction effects of c-fANOVA.
Individual Conditional Expectations (ICE).
$\phi^{\text{ICE}}(S):=F(b_{S},x_{-S})$
@@ E.1.2 Gradient-Based Attribution Methods
DeepSHAP (Lundberg and Lee,, 2017) computes attributions layer-wise, where the SV is used. It was shown by Deng et al., (2024), that for each layer, the layer-wise attribution is given by
$\phi(i)=\sum_{\kappa\in\Omega_{i}}\phi(\kappa)+\sum_{S\supset i:s>1}\frac{1}{s}\sum_{\kappa\in\Omega_{S}}I(\kappa)=\psi(i)+\sum_{S\supset i:s>1}\frac{1}{s}J(S)=f_{i}^{(b)}+\sum_{S\supset i:s>1}\frac{1}{s}f^{(b)}_{S},$
@@
i.e., applying IG for each feature. By taking the gradient, all effects vanish that do not involve feature $x_{i}$, as seen in Eq. (5). In detail, for IG, we have
@@ F.1 Synthetic Examples
To model the data-generating processes $F_{\text{lin}}$, $F_{\text{int}}$, and $F_{\text{add}}$, we randomly draw 10,000 samples and fit correctly specified linear models with sklearn (Pedregosa et al.,, 2011), ensuring correctly learned effects for each functional form. Input features are transformed to match the underlying functional relationships, with the relevant main effects and interaction terms selected accordingly. This method guarantees that the models retain linearity in the parameter space, consistent with the specified functional relationships. As a result, we obtain an almost perfect fit across all three synthetic datasets, achieving $R^{2}$ values greater than 0.99 on a separately sampled test dataset of size 10,000 following the same distribution as defined above.
@@ F.2 Models and Datasets for Real-World Examples
  • Random Forest: A random forest model with 50 trees was trained on the preprocessed data, achieving an $R^{2}$ of $0.9471\pm 0.0024$ and an MSE of 1 863±103plus-or-minus18631031\,863$\pm$ 1031 863 ± 103.
  • Neural Network: A fully connected neural network with four hidden layers (sizes 100, 50, 10, 1) was trained on the preprocessed data, where the categorical features were treated as numeric and standardized. The regression target variable was additionally log-normalized. This model achieved an $R^{2}$ of $0.9436\pm 0.0016$ and an MSE of $0.3299\pm 0.0055$.
@@
The Titanic dataset is used for a binary classification task that predicts the survival probability of passengers based on characteristics such as age, sex, class of travel, and fare. We preprocess the data by applying an OrdinalEncoder to the categorical features.
@@
The sentiment analysis language data is based on the IMDB dataset (Maas et al.,, 2011; Lhoest et al.,, 2021) where example sentences to the model are self-created. The model is a pre-trained and fine-tuned version of DistilBERT (Sanh et al.,, 2019) on IMDB as provided with the transformers (Wolf et al.,, 2020) framework. The model predicts the sentiment of movie reviews with a sentiment score $\left[-1,1\right]$ where values close to $-1$ denote a negative predicted sentiment and values close to $1$ correspond to a positive predicted sentiment. Sentiment scores close to 00 are considered neutral. In our experiments, we manually remove punctuation and special symbols. In case of b-fANOVA, we set out-of-coalition tokens to the MASK token of the transformer model. Notably, the selected DistilBERT variant assigns sentences provided only MASK tokens moderately high scores of around $0.50$ (i.e., on the positive side). We do not evaluate and report performance scores of the pre-trained model on a separate test set.